You are conducting a multinomial hypothesis test for the claim that the 4 categories occur with the following frequencies: H:Please explain by using the TI-83 calculator if possible.
Answer #1

Answer

total frequency n= 120

Expected frequency

Category Expected
A 0.25*120 = 30
B 0.4*120 = 48
C 0.1*120=12
D 0.25*120=30

Using TI 84 calculator

press STAT then EDIT and enter the obsered frequency and expected frequency in L1 and L2 lists.

Again

press STAT then TESTS and select X? GOF

select L1 as observed value

L2 as expected values

df= 3 (number of categories -1)

Press Calculate

we get

x2 = 5.267